OpenStudy (anonymous):

If f (x) = −2x 2 + 5x − 2, find f (−3)

OpenStudy (saifoo.khan):

37

OpenStudy (anonymous):

f(-3) = -2(-3)2 +5(-3) - 2 = -2 * 9 - 15 - 2 = - 18 -15 - 2 = -35